OpenStudy (anonymous):

Equate the imaginary parts of this equation: cos^2θ+2isinθcosθ-sin^2θ=cos2θ+isinθ , Please show steps. Thank you...

OpenStudy (anonymous):

do you know which are the imaginary parts?

OpenStudy (anonymous):

yes The parts with the I - that is 2isinθ and isinθ

OpenStudy (anonymous):

in the first one you are missing something out

OpenStudy (anonymous):

is that the whole term - 2isinθcosθ

OpenStudy (anonymous):

yep (but this "-" is minus or hyphen?)

OpenStudy (anonymous):

sorry forget the minus

OpenStudy (anonymous):

it is just 2isinθcosθ

OpenStudy (anonymous):

Now you can equate them. What is next? You need to solve for "theta"?

OpenStudy (anonymous):

Okay do I do transposition of formula?

OpenStudy (anonymous):

My problem is that I am stuck at this particular point.

OpenStudy (anonymous):

imaginary terms are \[2 \sin \theta \cos \theta \] \[\sin \theta\] now \[2 \sin \theta \cos \theta=\sin \theta\] so here \[\sin 2 \theta= \sin \theta\]

OpenStudy (anonymous):

theyatin thanks but my textbook gives a different answer, let me see what carlosgp says. :)

OpenStudy (anonymous):

you have:\[2\sin(\theta)\cos(\theta)=\sin(\theta)\] The trivial solution is theta=0 that will make sin(theta)=0 For another solution, suppose sin(theta)<>0, then cos(theta)=1/2, find the "theta" whose cosine is 0.5

OpenStudy (anonymous):

okay, the answer from the textbook is sin2θ = 2sinθcosθ

OpenStudy (anonymous):

now you will have to solve for value of theta and that is zero. . . beacause \[2 \theta = \theta\] its the only case if theta is zero

OpenStudy (anonymous):

thanks carlos, theyatin

OpenStudy (anonymous):

Any idea why the txt book gives this answer sin2θ = 2sinθcosθ

OpenStudy (anonymous):

No clue

OpenStudy (anonymous):

okay many thanks

OpenStudy (anonymous):

u sure its not printing mistake?? we did everything right. . .

OpenStudy (vincent-lyon.fr):

That's a typo, second part must be sin2θ

OpenStudy (anonymous):

@theyatin , the question specifically says to equate the imaginary parts. It didn't require solving, just equating. Kind regards....

OpenStudy (anonymous):

@unseenoceans oh goodness then you have your answer already right?? i over explained. . .